3(2h + 1) – 2(h – 4) = 21
1
Expert's answer
2013-03-29T12:01:57-0400
3(2h + 1) - 2(h - 4) = 21
6h + 3 - 2h + 8 = 21
6h - 2h = 21 - 3 - 8
4h = 10
h = 10/4
h = 2.5
So,
Answer: h=2.5